HIV Score: 30 56%, 183 .2.28 Queston Holp Amanufactrer knows that on average 20% of toasters pr deed probablity that at least x of them will require repairs is less than 0.48. (b) the probability that at least y of them well not require repairs is greater than 0 78 tequre repars witar, 1 year hr teyaro sold then 17 toasters aro randomly uictod, td the smalost nu ber x and ︸ umber y suchthat a na binomial probablity.sums ca) xu (Type a whole rurnber )
Answer #1

Solution:

Part a

We are given p=0.20, n=17

P(X≥x) = 0.48

P(X<x) = 1 – 0.48 = 0.52

So, by using Binomial table,

X = 3 approximately

Part b

We are given p = 1 – 0.20 = 0.80, n = 17

P(Y≥y) = 0.78

P(Y<y) = 1 – 0.78 = 0.22

So, by using binomial table,

Y = 12 approximately